# The tidal evolution of anisotropic subhaloes: A new pathway to creating isotropic and cored satellites

### v1

@@ 4.2 Tidal isotropisation
The top panel shows the results for our fiducial circular orbit with $\mathcal{R}_{E}=0.25$. Note that $\beta_{50\%}$ for the isotropic subhalo remains close to zero at all times, while that of the tangentially anisotropic orbit increases slightly from its initial value of $-0.5$ to $-0.3$, after which it remains nearly constant for the duration of the simulation. The situation for the OM and radially anisotropic subhaloes is very different. They all undergo rapid isotropisation, with $\beta_{50\%}$ dropping close to zero by $\sim 1\,{\rm Gyr}$, which is only one-third of the orbital period. The middle panel of Fig. 11 shows the temporal evolution of $\beta_{50\%}$ for our subhaloes along a highly eccentric orbit with $\mathcal{R}_{E}=0.25$ and $e=0.9$. Overall the results are similar to those for the circular orbit; once again the isotropic subhalo remains isotropic, while the OM and radially anisotropic subhaloes have $\beta_{50\%}$ drop rapidly to zero during the first Gyr of evolution. In the case of the tangentially anisotropic subhalo the evolution of $\beta_{50\%}$ is somewhat different. In particular, at late times the subhalo’s velocity anisotropy becomes more and more isotropic over time, in a step-wise fashion.
@@
We emphasise that we find tidal isotropisation to be present along all orbits that we have explored ($0.15\leq\mathcal{R}_{E}\leq 1.0$; $0\leq e\leq 0.9$), not just the two orbits depicted in Fig. 11. In general, the rate of tidal isotropisation is faster and more pronounced if the maximum tidal strength along the orbit is stronger. Hence, at fixed $\mathcal{R}_{E}$, isotropisation is faster along more radial orbits. Similarly at fixed orbital eccentricity, isotropisation is more rapid along more bound orbits (lower $\mathcal{R}_{E}$).
@@
What is the cause of this tidal isotropisation? Several mechanisms can be at play. First of all, as we have seen above, tides preferentially strip the more radial, low-angular momentum orbits, which have larger apo-centers than their high-angular momentum, more circular counterparts with the same orbital energy. Hence, stripping should naturally lead to a reduction of $\beta$. We believe that this is the main cause for the rapid isotropisation of the radially anisotropic subhaloes (but see below). Another mechanism that is likely to play a role is the re-virialisation that supersedes the mass stripping. During re-virialisation, particles change their energy and angular momentum in response to a time-varying potential (violent relaxation), and this tends to drive the system to isotropy, which corresponds to the lowest energy state (Doremus et al., 1971). This is likely to be the main mechanism causing the slow, step-wise isotropisation of the tangentially anisotropic subhalo along the highly eccentric orbit; each ‘step’ here corresponds to a peri-centric passage, during which the subhalo experiences an impulsive shock. Following each shock, the subhalo rapidly re-virialises, after which it is left with a value of $\beta_{50\%}$ that is somewhat closer to isotropy. An impulsive shock gives each particle in the subhalo a velocity-impulse, $\Delta{\bf v}({\bf x})$, which changes both the orbital energy and angular momentum of the particle. During the subsequent re-virialisation, both the energy and angular momentum of the particles once again undergo changes due to the time-varying potential. The net outcome of this repeated mixing in energy-angular momentum space will be to drive the system to a more isotropic state.
@@ 4.3 Breaking the mass–anisotropy degeneracy of MW satellites
In order to better quantify this effect, we run a suite of simulations that tidally evolve the $\beta=+0.5$ subhalo on orbits that densely sample the entire orbital parameter space, covering $0.1\leq\mathcal{R}_{E}\leq 1.2$ and $0\leq e\leq 1$. The colour-shading in Fig. 14 indicates the velocity anisotropy $\overline{\beta_{50\%}}$, time-averaged over $0.5t_{\text{orb}}$$t_{\text{orb}}$. Overall, the effect of tidal isotropisation becomes more pronounced with increasing maximal tidal field strength (equivalent to decreasing peri-centric distance), which corresponds to $e\rightarrow 1$ and/or $\mathcal{R}_{E}\rightarrow 0$), reducing the initial $\beta_{50\%}=0.5$ (black) down to anywhere between $\overline{\beta_{50\%}}\simeq 0.3$ (orange) and $0.0$ (blue) during first infall. The fact that the trend slightly reverses towards the bottom-right corner of the $e$$\mathcal{R}_{E}$ parameter space is simply due to the fact that along the highly radial orbits the tidal perturbations become highly impulsive, resulting in mass loss less than that along a circular orbit of similar $\mathcal{R}_{E}$.
The fact that all subhaloes experience rapid tidal isotropisation allows us to place conservative constraints on the individual velocity anisotropy of the MW dwarf spheroidal satellite galaxies, whose measured orbital parameters from Pace et al. 2022 are overlaid in Fig. 14 as gray and black error bars. Since subhaloes only become more and more isotropic over time, and many MW satellite galaxies have been orbiting the MW halo for more than a single orbit (Fillingham et al., 2019; Battaglia et al., 2022), the $\overline{\beta_{50\%}}(e,\mathcal{R}_{E})$ shown in Fig. 14 may be considered a conservative upper limit for their amount of radial anisotropy. In the case of Segue 1, Segue 2, Coma Berenices, Tucana IV, Draco II, Boötes II and Triangulum II we infer that $\overline{\beta_{50\%}}<0.1$. Hence, we may safely assume that at present these systems should be close to isotropic, removing any uncertainties regarding their inferred density profiles that result from the mass-anisotropy degeneracy. We emphasise, though, that we have assumed that the stellar body of a satellite galaxy undergoes the same level of tidal isotropisation as the subhalo it inhabits. To what extent this is accurate will need to be tested using idealised simulations of two-component anisotropic halo$+$star systems, which we leave for future work.
@@ 5 Summary and Conclusions
  • Radially (tangentially) anisotropic subhaloes experience more (less) tidal mass loss than their isotropic counterparts, irrespective of the subhalo’s orbital parameters. Quantitatively, the bound mass fraction of an anisotropic subhalo can already differ from that of the conventional, isotropic predictions by order unity after a single orbit and by several orders of magnitude after a Hubble time. This strong dependence of the tidal evolution on the internal orbital anisotropy is a consequence of the fact that more radial orbits have larger apo-centers compared to more tangential orbits of the same orbital energy. Since stripping mainly removes orbits that reach outside of the tidal radius, more radial orbits are therefore more easily stripped.
@@ Acknowledgements
We acknowledge useful conversations with Raphaël Errani, Marla Geha and Go Ogiya. FvdB is supported by the National Science Foundation (NSF) through grants AST-2307280 and AST-2407063. This research is partially supported by the National Science and Technology Council (NSTC) of Taiwan under Grant No. NSTC 111-2628-M-002-005-MY4 and the NTU Academic Research-Career Development Project under Grant No. NTU-CDP-113L7729. We have used yt (Turk et al., 2011) for data visualisation.
@@ Appendix A Numerical Convergence and Stability Tests
Figure 15: Time evolution of shell-averaged density (top row) and 3D velocity dispersion profiles (bottom) of the isotropic, $\beta=0.1,0.2,$ and $-0.5$ subhaloes (leftmost to rightmost columns) in isolation for 50 Gyr $\simeq 25t_{\text{cross}}$ with the AMR and simulation setup described in §2.2. Artificial two-body heating remains negligible down to grid resolution scale $r=0.012r_{\text{s,0}}$, demonstrating both the stability of PIANISTpy-constructed phase-space-truncated initial conditions and numerical accuracy of gamer-2 over the resolved range of length scales.
@@ A.1 Numerical stability and heating in isolation
To verify that our AMR refinement strategy (see §2.2) is adequate, and that our initial conditions are stable, we dynamically evolve the isotropic, $\beta=+0.2,+0.1,$ and $-0.5$ haloes, generated with $N_{\text{par}}=10^{7}$ equal-mass particles using PIANISTpy, in isolation for $50\,{\rm Gyr}\simeq 25t_{\text{cross}}$. Fig. 15 shows the resulting density (top row) and 3D velocity dispersion profiles (bottom) at four different epochs: $t=0$ (black dashed), 3 (blue), 14 (orange), and $50$ Gyr (red). Different columns correspond to haloes with a different initial anisotropy, as indicated. Evidently, the isotropic and $\beta=+0.1$ and $-0.5$ systems are extremely stable, showing no sign of evolution from the initial conditions for the full duration of each simulation. Hence, we see that, gamer-2, with the ARM refinement strategy adopted here, provides sufficient force and mass resolutions, free of any discernible numerical two-body relaxation effect. In the case of $\beta=+0.2$, some late-time evolution in the 3D velocity dispersion is apparent inside of $0.1r_{\text{s,0}}$. As discussed in detail in Chiang & van den Bosch (2024), this owes to the onset of the ROI (cf. Fig. 13).
@@ A.2 Code comparison
As this is the first time that gamer-2 is used to study the tidal evolution of substructure it is judicious to compare its performance with a more tested code. We therefore aim to reproduce some results of van den Bosch & Ogiya (2018), who used the hierarchical $N$-body code treecode, written by J. Barnes and with some improvements due to J. Dubinski, to evolve isotropic $N$-body subhaloes in the analytical potential of a host halo. treecode uses a Barnes & Hut (1986) octree to compute accelerations based on a multipole expansion up to quadrupole order, and uses a second-order leap-frog integration scheme to solve the equations of motion. Forces between particles are softened using a simple Plummer softening.
@@
We use gamer-2 to evolve the IC files of van den Bosch & Ogiya (2018) (hard-truncated, isotropic NFW subhaloes with $N_{\text{par}}=10^{5}$) along circular orbits in the same analytical host halo as used in their study. Fig. 16 compares the resulting bound mass fractions as function of time (dashed lines) to those of van den Bosch & Ogiya (2018) (solid lines, taken from their Fig. 4). Different colours correspond to different circular orbits, with $\mathcal{R}_{E}$ ranging from $1.0$ (blue) to $0.1$ (red) in steps of $0.1$. Note the excellent agreement between the gamer-2 and treecode results, with relative differences smaller than $|\Delta f_{\rm bound}(t)/f_{\rm bound}(t)|\leq 5\%$ across all ten cases simulated. This validates gamer-2 as a suitable and competitive $N$-body code.

### v2

@@ 4.2 Tidal isotropisation
The top panel shows the results for our fiducial circular orbit with $\mathcal{R}_{E}=0.25$. Note that $\beta_{50\%}$ for the isotropic subhalo remains close to zero at all times, while that of the tangentially anisotropic orbit increases slightly from its initial value of 0.5-0.5 to 0.3-0.3, after which it remains nearly constant for the duration of the simulation. The situation for the OM and radially anisotropic subhaloes is very different. They all undergo rapid isotropisation, with $\beta_{50\%}$ dropping close to zero by $\sim 1\,{\rm Gyr}$, which is only one-third of the orbital period. The middle panel of Fig. 11 shows the temporal evolution of $\beta_{50\%}$ for our subhaloes along a highly eccentric orbit with $\mathcal{R}_{E}=0.25$ and $e=0.9$. Overall the results are similar to those for the circular orbit; once again the isotropic subhalo remains isotropic, while the OM and radially anisotropic subhaloes have $\beta_{50\%}$ drop rapidly to zero during the first Gyr of evolution. In the case of the tangentially anisotropic subhalo the evolution of $\beta_{50\%}$ is somewhat different. In particular, at late times the subhalo’s velocity anisotropy becomes more and more isotropic over time, in a step-wise fashion.
@@
We emphasise that we find tidal isotropisation to be present along all orbits that we have explored ($0.15\leq\mathcal{R}_{E}\leq 1.0$; $0\leq e\leq 0.9$), not just the two orbits depicted in Fig. 11. We have also investigated the evolution of our subhaloes in the energy–angular momentum space, measured in their rest frames, which clearly reveals isotropisation in that the distribution of angular momentum at a given energy becomes more homogeneous with time. In general, the rate of tidal isotropisation is faster and more pronounced if the maximum tidal strength along the orbit is stronger. Hence, at fixed $\mathcal{R}_{E}$, isotropisation is faster along more radial orbits. Similarly at fixed orbital eccentricity, isotropisation is more rapid along more bound orbits (lower $\mathcal{R}_{E}$).
@@
What is the cause of this tidal isotropisation? Several mechanisms can be at play. First of all, as we have seen above, tides preferentially strip the more radial, low-angular momentum orbits, which have larger apo-centres than their high-angular momentum, more circular counterparts with the same orbital energy. Hence, stripping should naturally lead to a reduction of $\beta$. We believe that this is the main cause for the rapid isotropisation of the radially anisotropic subhaloes (but see below). Another mechanism that is likely to play a role is the re-virialisation that supersedes the mass stripping. During re-virialisation, particles change their energy and angular momentum in response to a time-varying potential (violent relaxation), and this tends to drive the system to isotropy, which corresponds to the lowest energy state (Doremus et al., 1971). This is likely to be the main mechanism causing the slow, step-wise isotropisation of the tangentially anisotropic subhalo along the highly eccentric orbit; each ‘step’ here corresponds to a peri-centric passage, during which the subhalo experiences an impulsive shock. Following each shock, the subhalo rapidly re-virialises, after which it is left with a value of $\beta_{50\%}$ that is somewhat closer to isotropy. An impulsive shock gives each particle in the subhalo a velocity-impulse, $\Delta{\bf v}({\bf x})$, which changes both the orbital energy and angular momentum of the particle. During the subsequent re-virialisation, both the energy and angular momentum of the particles once again undergo changes due to the time-varying potential. The net outcome of this repeated mixing in energy-angular momentum space will be to drive the system to a more isotropic state.
@@ 4.3 Breaking the mass–anisotropy degeneracy of MW satellites
[…]
The fact that all subhaloes experience rapid tidal isotropisation allows us to place conservative constraints on the individual velocity anisotropy of the MW dwarf spheroidal satellite galaxies, whose measured orbital parameters from Pace et al. 2022 are overlaid in Fig. 14 as gray and black error bars. Since subhaloes only become more and more isotropic over time, and many MW satellite galaxies have been orbiting the MW halo for more than a single orbit (Fillingham et al., 2019; Battaglia et al., 2022), the $\overline{\beta_{50\%}}(e,\mathcal{R}_{E})$ shown in Fig. 14 may be considered a conservative upper limit for their amount of radial anisotropy. In the case of Segue 1, Segue 2, Coma Berenices, Tucana IV, Draco II, Boötes II and Triangulum II we infer that $\overline{\beta_{50\%}}<0.1$. Hence, we may safely assume that at present these systems should be close to isotropic, removing any uncertainties regarding their inferred density profiles that result from the mass-anisotropy degeneracy. We emphasise, though, that we have assumed that the stellar body of a satellite galaxy undergoes the same level of tidal isotropisation as the subhalo it inhabits. To what extent this is accurate will need to be tested using idealised simulations of two-component anisotropic halo++star systems, which we leave for future work.
@@ 5 Summary and Conclusions
  • Radially (tangentially) anisotropic subhaloes experience more (less) tidal mass loss than their isotropic counterparts, irrespective of the subhalo’s orbital parameters. Quantitatively, the bound mass fraction of an anisotropic subhalo can already differ from that of the conventional, isotropic predictions by order unity after a single orbit and by several orders of magnitude after a Hubble time. This strong dependence of the tidal evolution on the internal orbital anisotropy is a consequence of the fact that more radial orbits have larger apo-centres compared to more tangential orbits of the same orbital energy. Since stripping mainly removes orbits that reach outside of the tidal radius, more radial orbits are therefore more easily stripped.
@@ Acknowledgements
We acknowledge useful conversations with Raphaël Errani, Marla Geha and Go Ogiya. We also thank the anonymous referee for a constructive and detailed report. FvdB is supported by the National Science Foundation (NSF) through grants AST-2307280 and AST-2407063. This research is partially supported by the National Science and Technology Council (NSTC) of Taiwan under Grant No. NSTC 111-2628-M-002-005-MY4 and the National Taiwan University (NTU) Academic Research-Career Development Project under Grant No. NTU-CDP-113L7729. We use NumPy (Harris et al., 2020) and SciPy (Virtanen et al., 2020) for data analysis, and Matplotlib (Hunter, 2007) and yt (Turk et al., 2011) for data visualisation.
@@ Appendix A Numerical Convergence and Stability Tests
Figure 15: Time evolution of shell-averaged density (top row) and 3D velocity dispersion profiles (bottom) of the isotropic, $\beta=0.1,0.2,$ and 0.5-0.5 subhaloes (leftmost to rightmost columns) in isolation for 50 Gyr $\simeq 25t_{\text{cross}}$ with the AMR and simulation setup described in §$2.$ Artificial two-body heating remains negligible down to grid resolution scale $r=0.012r_{\text{s,0}}$, demonstrating both the stability of PIANISTpy-constructed phase-space-truncated initial conditions and numerical accuracy of gamer-2 over the resolved range of length scales.
@@ A.1 Numerical stability and heating in isolation
To verify that our AMR refinement strategy (see §2.2) is adequate, and that our initial conditions are stable, we dynamically evolve the isotropic, $\beta=+0.2,+0.1,$ and 0.5-0.5 haloes, generated with $N_{\text{par}}=10^{7}$ equal-mass particles using PIANISTpy, in isolation for $50\,{\rm Gyr}\simeq 25t_{\text{cross}}$. Fig. 15 shows the resulting density (top row) and 3D velocity dispersion profiles (bottom) at four different epochs: $t=0$ (black dashed), 3 (blue), 14 (orange), and $50$ Gyr (red). Different columns correspond to haloes with a different initial anisotropy, as indicated. Evidently, the isotropic and $\beta=+0.1$ and 0.5-0.5 systems are extremely stable, showing no sign of evolution from the initial conditions for the full duration of each simulation. Hence, we see that, gamer-2, with the ARM refinement strategy adopted here, provides sufficient force and mass resolutions, free of any discernible numerical two-body relaxation effect. In the case of $\beta=+0.2$, some late-time evolution in the 3D velocity dispersion is apparent inside of $0.1r_{\text{s,0}}$. As discussed in detail in Chiang et al. (in preparation), this owes to the onset of the ROI (cf. Fig. 13).
@@ A.2 Code comparison
As this is the first time that gamer-2 is used to study the tidal evolution of substructure, it is judicious to compare its performance with a more tested code. We therefore aim to reproduce some results of van den Bosch & Ogiya (2018), who used the hierarchical $N$-body code treecode, written by J. Barnes and with some improvements due to J. Dubinski, to evolve isotropic $N$-body subhaloes in the analytical potential of a host halo. treecode uses a Barnes & Hut (1986) octree to compute accelerations based on a multipole expansion up to quadrupole order, and uses a second-order leap-frog integration scheme to solve the equations of motion. Forces between particles are softened using a simple Plummer softening.
@@
We use gamer-2 to evolve the initial condition files of van den Bosch & Ogiya (2018) (hard-truncated, isotropic NFW subhaloes with $N_{\text{par}}=10^{5}$) along circular orbits in the same analytical host halo as used in their study. Fig. 16 compares the resulting bound mass fractions as function of time (dashed lines) to those of van den Bosch & Ogiya (2018) (solid lines, taken from their Fig. 4). Different colours correspond to different circular orbits, with $\mathcal{R}_{E}$ ranging from $1.0$ (blue) to $0.1$ (red) in steps of $0.1$. Note the excellent agreement between the gamer-2 and treecode results, with relative differences smaller than $|\Delta f_{\rm bound}(t)/f_{\rm bound}(t)|\leq 5\%$ across all ten cases simulated. This validates gamer-2 as a suitable and competitive $N$-body code.